# Exploiting Agent Symmetries for Performance Analysis of Distributed Optimization Methods

### v1

@@ 3 Equivalence classes of agents in PEP
Let $F$ and $G=P^{T}P$ be a solution of an agent-dependent PEP, described in Section 2. We can order their elements to regroup them depending on the agent to which they are associated:
$F=\begin{bmatrix}f_{1}^{T}&\dots&f_{n}^{T}\end{bmatrix},\quad\text{where $f_{i}=\quantity[f_{i}^{k}]_{k\in I}\in\mathbb{R}^{q}$ is a vector with the $q$ functions values of agent $i$,}$ (69)
@@
$P=\begin{bmatrix}P_{1}&\dots&P_{n}\end{bmatrix}\quad\text{where $P_{i}\in\mathbb{R}^{d\times p}$ contains the $p$ vector variables related to with agent $i$,}$ (70)
e.g. $P_{i}=\quantity[y_{i}^{k}\leavevmode\nobreak\ x_{i}^{k}\leavevmode\nobreak\ g_{i}^{k}]_{k\in I}$, and the Gram matrix $G\in\mathbb{R}^{np\times np}$ is thus defined as
@@
Therefore, diagonal block $G_{ii}$ are symmetric and off-diagonal blocks are such that $G_{ij}=G_{ji}^{T}$. We assume that each block $P_{i}$ (and $f_{i}$) contains the same type and number of variables for each agent $i$. The variables common to all agents, such as $x^{*}$, are copied into each agent block $P_{i}$. This also covers the case where agents are heterogeneous and hold different types or numbers of variables since we can always add variables as columns of $P_{i}$ (and $f_{i}$), even if agent $i$ does not use them. Moreover, each agent block $P_{i}$ (and $f_{i}$) has the same column order, so we can work with the same coefficient vectors for each agent. A coefficient vector for a variable $x$ is denoted $e_{x}\in\mathbb{R}^{p}$, and contains linear coefficients selecting the correct combination of columns in $P_{i}\in\mathbb{R}^{d\times p}$ to obtain vector $x_{i}\in\mathbb{R}^{d}$, i.e. $P_{i}e_{x}=x_{i}$, for any $i=1,\dots,n$. This notation allows, for example, to write $x_{i}^{T}x_{i}$ as $x_{i}^{T}x_{i}=e_{x}^{T}P_{i}^{T}P_{i}e_{x}=e_{x}^{T}G_{ii}e_{x}$, for any agent $i$. Similarly, let $e_{f(x)}$ be a vector of coefficients selecting the correct element in the vector $f_{i}$ such that $e_{f(x)}^{T}f_{i}=f_{i}(x_{i})$. These coefficient vectors will be used to explicit our new PEP formulations in Sections 4 and 5, as well as in Appendix A.
We use this agent blocks notation for $F$ and $G$ to define an equivalence relation between agents in Performance Estimation Problems (PEP).
@@
Let $(F,G)$ be a feasible solution of an agent-dependent PEP formulation for distributed optimization (14), written in the form of (69) and (71), and $(F_{\Pi},G_{\Pi})$ be another solution, obtained after permutation of agents $i,j\in\mathcal{V}$ in $F,G$. We define an equivalence relation between the two agents $i\sim j$ if both solutions $(F,G)$ and $(F_{\Pi},G_{\Pi})$ are feasible for the PEP and provide the same performance value,
$\mathcal{P}(F,G)=\mathcal{P}(F_{\Pi},G_{\Pi}).$
@@
Let $\mathcal{V}$ be a set of $n$ agents. The equivalence relation from Definition 8 implies a partition $\mathcal{T}$ of set $\mathcal{V}$ into $m$ equivalence classes (or subsets) of agents: $\mathcal{T}=\{\mathcal{V}_{1},\dots,\mathcal{V}_{m}\}$. Each class only contains agents that are equivalent to each other. The size of each class $\mathcal{V}_{u}$ is denoted $n_{u}$. We also denote the equivalence class of a given agent $i$ by $\mathcal{V}_{u_{i}}$.
Proposition 6 proves the existence of PEP solution for which equivalent agents have equal blocks in the solution. This will be the building block of our compact PEP formulations in Sections 4 and 5.
Proposition 6 (Existence of symmetric solution of a PEP).
Let $F\in\mathbb{R}^{nq}$ and $G\in\mathbb{R}^{np\times np}$ be any feasible solution of an agent-dependent PEP formulation for distributed optimization (14), and $\mathcal{T}=\{\mathcal{V}_{1},\dots,\mathcal{V}_{m}\}$ be the partition of $\mathcal{V}$ induced by the equivalence relation from Definition 8. There is a symmetrized agent-class solution $(F^{s}_{\mathcal{T}},G^{s}_{\mathcal{T}})$, with equal blocks for equivalent agents, providing another valid solution for the PEP,
[…]
Moreover, this symmetrized solution $(F^{s}_{\mathcal{T}},G^{s}_{\mathcal{T}})$ provides the same performance value as $(F,G)$
$\mathcal{P}(F^{s}_{\mathcal{T}},G^{s}_{\mathcal{T}})=\mathcal{P}(F,G).$
@@
By definition, any permutation of equivalent agents $\Pi\in\mathbb{R}^{n\times n}$ applied on a given PEP solution $F$, $G$ provides another valid PEP solution, with the same objective value:
$F_{\Pi}=F\quantity(\Pi\otimes I_{q}),\qquad P_{\Pi}=P\quantity(\Pi\otimes I_{p}),\qquad G_{\Pi}=P_{\Pi}^{T}P_{\Pi}=\quantity(\Pi\otimes I_{p})^{T}G\quantity(\Pi\otimes I_{p}).$
@@
A permutation of agents corresponds to a permutation of (sets of) columns and then the permutation matrix multiplies $F$ and $P$ on the right. The Kronecker products $\otimes$ are used to permute all the variables related to the same agent at the same time.
@@
Since the PEP problem is linear in $F$ and $G$, the combination of PEP solutions with the same objective value will give another PEP solution with the same objective value. We can thus construct a symmetrized PEP solution with the same objective value as $(F,G)$ by averaging all the possible permuted solutions $(F_{\Pi},G_{\Pi})$. Each permutation only involves agents from the same equivalence class, and therefore, the average applies to each class separately:
$f_{i}^{s}$ $=\frac{(n_{u_{i}}-1)!}{n_{u_{i}}!}\sum_{k\in\mathcal{V}_{u_{i}}}f_{k}=\frac{1}{n_{u_{i}}}\sum_{k\in\mathcal{V}_{u_{i}}}f_{k},$ for all $\text{for all }i\in\mathcal{V},$
$G_{ii}^{s}$ $=\frac{(n_{u_{i}}-1)!}{n_{u_{i}}!}\sum_{k\in\mathcal{V}_{u_{i}}}G_{kk}=\frac{1}{n_{u_{i}}}\sum_{k\in\mathcal{V}_{u_{i}}}G_{kk}$ for all $\text{for all }i\in\mathcal{V},$
[…]
Remark.
@@
The blocks composing $F^{s}$ (72) are denoted by $f_{A}^{u}$. For the symmetrized Gram matrix $G^{s}$ (76), we denote the three types of blocks composing it by $G_{A}^{u}$, $G_{B}^{u}$, and $G_{E}^{uv}$, for all equivalence classes $u,v=1,\dots,m$.
@@
  • Blocks $G_{A}^{u}\in\mathbb{R}^{p\times p}$ (for $u=1,\dots,m$) correspond to the blocks of the Gram matrix containing the scalar products between the variables of one agent from $\mathcal{V}_{u}$. These blocks lie on the diagonal of $G^{s}$ and are symmetric by definition.
@@
  • Blocks $G_{B}^{u}\in\mathbb{R}^{p\times p}$ (for $u=1,\dots,m$) correspond to the blocks of the Gram matrix containing the scalar products between variables of two different agents from the same class $\mathcal{V}_{u}$. These blocks are symmetric because they can be written as a sum of symmetric matrices $\frac{1}{n_{u_{i}}(n_{u_{i}}-1)}\sum_{k,l\in\mathcal{V}_{u_{i}},k>l}(G_{kl}+G_{lk}^{T})$. Moreover, they are only defined for classes with at least 2 agents ($n_{u}\geq 2$).
@@
  • Blocks $G_{E}^{uv}\in\mathbb{R}^{p\times p}$ (for $u\neq v=1,\dots,m$) correspond to the blocks of the Gram matrix containing the scalar products between variables of two agents from different classes $\mathcal{V}_{u}$ and $\mathcal{V}_{v}$. These blocks are non-symmetric and are only defined when there are at least 2 different classes of agents ($m\geq 2$).
@@
Therefore, we can solve the agent-dependent PEP problem, described in Section 2, restricted to symmetric solutions of the form of Proposition 6, without impacting the resulting worst-case value. The symmetry in the solutions allows working with a limited number of variables $f_{A}^{u}$, $G_{A}^{u}$, $G_{B}^{u}$, and $G_{E}^{uv}$ in PEP, depending only on the (smaller) number of equivalent classes of agents $m$. This requires reformulating all the elements of the agent-dependent PEP (objective and constraints) in terms of these variables. Section 4 shows how to perform such a reformulation when all the agents are equivalent in the PEP, which occurs in many usual settings. This leads to a PEP formulation with only one equivalence class and hence whose size is independent of the number of agents. Section 5 generalizes reformulation techniques from Section 4 to a general number of equivalence classes.
@@ 4 Agent-independent PEP formulation when all agents are equivalent
All the agents are equivalent in the Performance Estimation Problem applied to a distributed algorithm, in the sense of Definition 8. Therefore, there is only one equivalence class which is $\mathcal{V}$. In other words, this assumption means that all the agents can be permuted in a PEP solution without impacting its worst-case value and thus no agent plays a specific role in the algorithm or its performance evaluation.
This assumption is satisfied in many usual performance evaluation settings $\mathcal{S}_{n}$, for which all agents play an identical role in the algorithm and its performance evaluation. Under Assumption 1, the symmetrized solution from Proposition 6 only has a few different blocks that repeat.
@@
When all the agents ($n\geq 2$) are equivalent, see Assumption 1, any PEP solution $(F,G)$ can be fully symmetrized, without impacting its worst-case value:
$F^{s}$ $=\quantity[f_{A}^{T}\dots f_{A}^{T}]$ with $\text{with }\leavevmode\nobreak\ f_{A}$ $=\frac{1}{n}\sum_{i=1}^{n}f_{i}\leavevmode\nobreak\ \leavevmode\nobreak\ \in\mathbb{R}^{q},$ (77)
[…]
@@
The results follow from applying Proposition 6 with $\mathcal{V}$ as the only equivalence class. Since there is only one equivalence class of agents, the Gram matrix only has two types of blocks, as explained in Remark Remark. ∎
We consider the agent-dependent PEP formulation from Section 2, with a given class of averaging matrix, and we constrain its solution to be fully symmetric, as expressed in (77) and (81), to obtain a compact PEP formulation with size independent of the number of agents. In a second time, we describe the implications of a fully symmetrized PEP solution on the worst-case local iterates and functions.

### v2

@@ 3 Equivalence classes of agents in PEP
Let $\mathbf{f}$ and $G=P^{T}P$ be a solution of an agent-dependent PEP, described in Section 2. We can order their elements to regroup them depending on the agent to which they are associated :
[…]
@@
[…]
e.g. $P_{i}=\quantity[y_{i}^{k}\penalty 10000\ x_{i}^{k}\penalty 10000\ g_{i}^{k}]_{k\in I}$, and the Gram matrix $G\in\mathbb{R}^{np\times np}$ is thus defined as
@@
Therefore, diagonal blocks $G_{ii}$ are symmetric and off-diagonal blocks are such that $G_{ij}=G_{ji}^{T}$. We assume that each block $P_{i}$ (and $f_{i}$) contains the same type and number of variables for each agent $i$. The variables common to all agents, such as $x^{*}$, are copied into each agent block $P_{i}$. This also covers the case where agents are heterogeneous and hold different types or numbers of variables since we can always add variables as columns of $P_{i}$ (and $f_{i}$), even if agent $i$ does not use them. Moreover, each agent block $P_{i}$ (and $f_{i}$) has the same column order, enabling us to use the same coefficient vectors for column selection in each agent block.
Notation (Coefficient vector $e_{x}\in\mathbb{R}^{p}$).
A coefficient vector for a variable $x$ is denoted $e_{x}\in\mathbb{R}^{p}$, and contains linear coefficients selecting the correct combination of columns in $P_{i}\in\mathbb{R}^{d\times p}$ to obtain vector $x_{i}\in\mathbb{R}^{d}$, i.e. $P_{i}e_{x}=x_{i}$, for any $i=1,\dots,n$. This notation allows, for example, to write $x_{i}^{T}x_{i}$ as $x_{i}^{T}x_{i}=e_{x}^{T}P_{i}^{T}P_{i}e_{x}=e_{x}^{T}G_{ii}e_{x}$, for any agent $i$. Similarly, let $e_{f(x)}$ be a vector of coefficients selecting the correct element in the vector $f_{i}$ such that $e_{f(x)}^{T}f_{i}=f_{i}(x_{i})$. These coefficient vectors will be used to explicit our new PEP formulations in Sections 4 and 5, as well as in Appendix A.
Using this agent block notation for $\mathbf{f}$ and $G$, we can define permutations of agents in a PEP solution, leading to the definition of an equivalence relation between agents in performance estimation problems (Definition 8 below).
Notation (Permuted solutions).
For any feasible solution $(\mathbf{f},G)$ of an agent-dependent PEP formulation (15), written in the form of (69) and (71), the solution obtained after permutation of agents $i,j\in\mathcal{V}$ in $\mathbf{f}$ and $G$ is denoted $(\mathbf{f}_{\Pi_{ij}},G_{\Pi_{ij}})$.
Example (Permuted solutions).
Let $(\mathbf{f},G)$ be a feasible PEP solution written as
$\hskip 49.79231pt\mathbf{f}$ $=\begin{bmatrix}f_{1}^{T}&f_{2}^{T}&\dots&f_{n}^{T}\end{bmatrix}$ $G$ $=P^{T}P\qquad\text{with }\quad P=\begin{bmatrix}P_{1}&P_{2}&\dots&P_{n}\end{bmatrix}$. (72)
Agents 11 and $2$ can be swapped to obtain a permuted solution $(\mathbf{f}_{\Pi_{12}},G_{\Pi_{12}})$ :
$\hskip 48.36958pt\mathbf{f}_{\Pi_{12}}$ $=\begin{bmatrix}f_{2}^{T}&f_{1}^{T}&\dots&f_{n}^{T}\end{bmatrix}$ $G_{\Pi_{12}}$ $=P_{\Pi_{12}}^{T}P_{\Pi_{12}}\qquad\text{with}\quad P_{\Pi_{12}}$ $=\begin{bmatrix}P_{2}&P_{1}&\dots&P_{n}\end{bmatrix}$. (73)
@@
Let $i,j\in\mathcal{V}$ be two distinct agents. We define an equivalence relation between the two agents $i\sim j$ if, for any feasible solution $(\mathbf{f},G)$ (69)-(71) of an agent-dependent PEP formulation (15), the permuted solution $(\mathbf{f}_{\Pi_{ij}},G_{\Pi_{ij}})$, with agent $i$ and $j$ swapped, is feasible for the PEP and provides the same performance value,
$\mathcal{P}(\mathbf{f},G)=\mathcal{P}(\mathbf{f}_{\Pi_{ij}},G_{\Pi_{ij}}).$
@@
Let $\mathcal{V}$ be a set of $n$ agents. The equivalence relation from Definition 8 implies a partition $\mathcal{T}$ of set $\mathcal{V}$ into $r$ equivalence classes (or subsets) of agents : $\mathcal{T}=\{\mathcal{V}_{1},\dots,\mathcal{V}_{r}\}$. Each class only contains agents that are equivalent to each other. The size of each class $\mathcal{V}_{u}$ is denoted $n_{u}$ ($u=1,\dots,r$). We also denote the equivalence class of a given agent $i$ by $\mathcal{V}_{u_{i}}$.
Proposition 6 proves the existence of PEP solutions for which equivalent agents have equal blocks in the solution. This will be the building block of our compact PEP formulations in Sections 4 and 5.
Proposition 6 (Existence of symmetric solutions in PEP).
Let $\mathbf{f}\in\mathbb{R}^{nq}$ and $G\in\mathbb{R}^{np\times np}$ be any feasible solution of an agent-dependent (and LMI Gram-representable) PEP formulation for distributed optimization (15) ; and $\mathcal{T}=\{\mathcal{V}_{1},\dots,\mathcal{V}_{r}\}$ be the partition of $\mathcal{V}$ induced by the equivalence relation from Definition 8. There is a symmetrized agent-class solution $(\mathbf{f}^{s}_{\mathcal{T}},G^{s}_{\mathcal{T}})$, with equal blocks for equivalent agents, providing another valid solution for the PEP,
$\hskip-11.38109pt\mathbf{f}^{s}_{\mathcal{T}}$ $=\quantity[(f_{1}^{s})^{T}\dots\,(f_{n}^{s})^{T}]\qquad\text{with }$ $f_{i}^{s}=f_{A}^{u_{i}}\vcentcolon=\frac{1}{n_{u_{i}}}\sum_{k\in\mathcal{V}_{u_{i}}}f_{k},\hskip 99.58464pt\text{for all }i\in\mathcal{V},$ (74)
$\hskip-11.38109ptG^{s}_{\mathcal{T}}$ $=\begin{bmatrix}G_{11}^{s}&G_{12}^{s}&\dots\\ G_{21}^{s}&\ddots&\\ \vdots&&G_{nn}^{s}\end{bmatrix}\quad\text{with }$ $G_{ij}^{s}=\begin{cases}G_{A}^{u_{i}}\vcentcolon=\frac{1}{n_{u_{i}}}\sum_{k\in\mathcal{V}_{u_{i}}}G_{kk}&\text{for $i=j$,}\\ G_{B}^{u_{i}}\vcentcolon=\frac{1}{n_{u_{i}}(n_{u_{i}}-1)}\sum_{k\in\mathcal{V}_{u_{i}}}\sum_{\begin{subarray}{c}l\in\mathcal{V}_{u_{i}}\\ l\neq k\end{subarray}}G_{kl}&\text{for $j\neq i,\penalty 10000\ j\in\mathcal{V}_{u_{i}}$},\\ G_{E}^{u_{i}u_{j}}\vcentcolon=\frac{1}{n_{u_{i}}}\sum_{k\in\mathcal{V}_{u_{i}}}G_{kj},&\text{for $j\neq i,\penalty 10000\ j\in\mathcal{V}\setminus\mathcal{V}_{u_{i}}$},\end{cases}$ (75)
Moreover, this symmetrized solution $(\mathbf{f}^{s}_{\mathcal{T}},G^{s}_{\mathcal{T}})$ provides the same performance value as $(\mathbf{f},G)$
$\mathcal{P}(\mathbf{f}^{s}_{\mathcal{T}},G^{s}_{\mathcal{T}})=\mathcal{P}(\mathbf{f},G).$
@@
By definition, any permutation of equivalent agents $\Pi\in\mathbb{R}^{n\times n}$ applied on a given PEP solution $\mathbf{f}$, $G$ provides another valid PEP solution, with the same objective value :
$\mathbf{f}_{\Pi}=\mathbf{f}\quantity(\Pi\otimes I_{q}),\qquad P_{\Pi}=P\quantity(\Pi\otimes I_{p}),\qquad G_{\Pi}=P_{\Pi}^{T}P_{\Pi}=\quantity(\Pi\otimes I_{p})^{T}G\quantity(\Pi\otimes I_{p}).$
@@
A permutation of agents corresponds to a permutation of (sets of) columns and then the permutation matrix multiplies $\mathbf{f}$ and $P$ on the right. The Kronecker products $\otimes$ are used to permute all the variables related to the same agent at the same time.
@@
Since the PEP problem is LMI Gram-representable, its objective is linear in $\mathbf{f}$ and $G$ and the combination of PEP solutions with the same objective value gives another PEP solution with the same objective value. We can thus construct a symmetrized PEP solution with the same objective value as $(\mathbf{f},G)$ by averaging all the possible permuted solutions $(\mathbf{f}_{\Pi},G_{\Pi})$. Each permutation exchanges only agents from the same equivalence class, and therefore, the average applies to each class separately :
$f_{i}^{s}$ $=\frac{(n_{u_{i}}-1)!}{n_{u_{i}}!}\sum_{k\in\mathcal{V}_{u_{i}}}f_{k}=\frac{1}{n_{u_{i}}}\sum_{k\in\mathcal{V}_{u_{i}}}f_{k}\penalty 10000\ \vcentcolon=f_{A}^{u_{i}},$ $\text{for all }i\in\mathcal{V},$
$G_{ii}^{s}$ $=\frac{(n_{u_{i}}-1)!}{n_{u_{i}}!}\sum_{k\in\mathcal{V}_{u_{i}}}G_{kk}=\frac{1}{n_{u_{i}}}\sum_{k\in\mathcal{V}_{u_{i}}}G_{kk}\penalty 10000\ \vcentcolon=G_{A}^{u_{i}}$ $\text{for all }i\in\mathcal{V},$
$G_{ij}^{s}$ $=\frac{(n_{u_{i}}-2)!}{n_{u_{i}}!}\sum_{k\in\mathcal{V}_{u_{i}}}\sum_{\begin{subarray}{c}l\in\mathcal{V}_{u_{i}}\\ l\neq k\end{subarray}}G_{kl}=\frac{1}{n_{u_{i}}(n_{u_{i}}-1)}\sum_{k\in\mathcal{V}_{u_{i}}}\sum_{\begin{subarray}{c}l\in\mathcal{V}_{u_{i}}\\ l\neq k\end{subarray}}G_{kl}\penalty 10000\ \vcentcolon=G_{B}^{u_{i}}$ $\text{for all $i\in\mathcal{V}$ and $j\neq i,\penalty 10000\ j\in\mathcal{V}_{u_{i}}$},$
$G_{ij}^{s}$ $=\frac{(n_{u_{i}}-1)!}{n_{u_{i}}!}\sum_{k\in\mathcal{V}_{u_{i}}}G_{ik}=\frac{1}{n_{u_{i}}}\sum_{k\in\mathcal{V}_{u_{i}}}G_{kj}\penalty 10000\ \vcentcolon=G_{E}^{u_{i}u_{j}}$ $\text{for all $i\in\mathcal{V}$ and $j\neq i,\penalty 10000\ j\in\mathcal{V}\setminus\mathcal{V}_{u_{i}}$},$
Definition 10 (Blocks $f_{A}^{u}$, $G_{A}^{u}$, $G_{B}^{u}$ and $G_{E}^{uv}$).
@@
The blocks repeating in $\mathbf{f}^{s}$ (74) are denoted by $f_{A}^{u}$, for each equivalence class $u=1,\dots,r$. For the symmetrized Gram matrix $G^{s}$ (75), we denote the three types of blocks composing it by $G_{A}^{u}$, $G_{B}^{u}$, and $G_{E}^{uv}$, for all equivalence classes $u,v=1,\dots,r$.
@@
  • Blocks $G_{A}^{u}\in\mathbb{R}^{p\times p}$ (for $u=1,\dots,r$) correspond to the blocks of the Gram matrix containing the scalar products between the variables of one agent from $\mathcal{V}_{u}$. These blocks lie on the diagonal of $G^{s}$ (75) and are symmetric by definition.
@@
  • Blocks $G_{B}^{u}\in\mathbb{R}^{p\times p}$ (for $u=1,\dots,r$) correspond to the blocks of the Gram matrix containing the scalar products between variables of two different agents from the same class $\mathcal{V}_{u}$. These blocks are symmetric because they can be written as a sum of symmetric matrices $G_{B}^{u}=\frac{1}{n_{u}(n_{u}-1)}\sum_{k,l\in\mathcal{V}_{u},k>l}(G_{kl}+G_{lk}^{T})$. Moreover, they are only defined for classes with at least 2 agents ($n_{u}\geq 2$).
@@
  • Blocks $G_{E}^{uv}\in\mathbb{R}^{p\times p}$ (for $u\neq v=1,\dots,r$) correspond to the blocks of the Gram matrix containing the scalar products between variables of two agents from different classes $\mathcal{V}_{u}$ and $\mathcal{V}_{v}$. These blocks are non-symmetric and are only defined when there are at least 2 different classes of agents ($r\geq 2$).
@@
Therefore, we can solve the agent-dependent PEP problem, described in Section 2, restricted to symmetric solutions of the form of Proposition 6, without impacting the resulting worst-case value. The symmetry in the solutions allows working with a limited number of variables $f_{A}^{u}$, $G_{A}^{u}$, $G_{B}^{u}$, and $G_{E}^{uv}$ in PEP, depending only on the (smaller) number of equivalent classes of agents $r$. This requires reformulating all the elements of the agent-dependent PEP (objective and constraints) in terms of these variables. Section 4 shows how to perform such a reformulation when all the agents are equivalent in the PEP, which occurs in many usual settings. This leads to a PEP formulation with only one equivalence class and hence whose size is independent of the number of agents. Section 5 generalizes reformulation techniques from Section 4 to a general number of equivalence classes.
@@ 4 Agent-independent PEP formulation when all agents are equivalent
All the agents are equivalent in the Performance Estimation Problem applied to a distributed algorithm, in the sense of Definition 8. Therefore, there is only one equivalence class which is $\mathcal{V}$.
In other words, this assumption means that all the agents can be permuted in a PEP solution without impacting its worst-case value and thus no agent plays a specific role in the algorithm or its performance evaluation. This assumption is satisfied in many usual performance evaluation settings $\mathcal{S}_{n}$, for which all agents play an identical role in the algorithm and its performance evaluation. Under Assumption 1, the symmetrized solution from Proposition 6 only has a few different blocks that repeat.
@@
When all the agents ($n\geq 2$) are equivalent, see Assumption 1, any solution $(\mathbf{f},G)$ of an agent-dependent (and LMI Gram-representable) PEP formulation for distributed optimization (15) can be fully symmetrized, without impacting its worst-case value :
$\mathbf{f}^{s}$ $=\quantity[f_{A}^{T}\dots f_{A}^{T}]$ $\text{with }\penalty 10000\ f_{A}$ $=\frac{1}{n}\sum_{i=1}^{n}f_{i}\penalty 10000\ \penalty 10000\ \in\mathbb{R}^{q},$ (76)
$G^{s}$ $=\begin{bmatrix}G_{A}&G_{B}&\dots\\ G_{B}&\ddots&\\ \vdots&&G_{A}\end{bmatrix}$ $\text{with }\penalty 10000\ G_{A}$ $=\frac{1}{n}\sum_{i=1}^{n}G_{ii}\penalty 10000\ \penalty 10000\ \in\mathbb{R}^{p\times p},$ $\hskip-8.53581pt\text{and}\penalty 10000\ G_{B}$ $=\frac{1}{n(n-1)}\sum_{i=1}^{n}\sum_{j\neq i}^{n}G_{ij}\penalty 10000\ \in\mathbb{R}^{p\times p}.$ (77)
@@
The results follow from applying Proposition 6 with $\mathcal{V}$ as the only equivalence class. Since there is only one equivalence class of agents, the Gram matrix only has two types of blocks, as explained in Definition 10. ∎
In this section, we consider the agent-dependent PEP formulation from Section 2, with a given class of averaging matrix, and we constrain its solution to be fully symmetric, as expressed in (76) and (77). First, we explain how and when this leads to a compact PEP formulation with size independent of the number of agents. Secondly, we describe the implications of a fully symmetrized PEP solution on the worst-case local iterates and functions.